OpenStudy (anonymous):

determine weather the graphs of the two equationare perpendicular, y 8x-5, 6y 5-x

OpenStudy (anonymous):

are the two equations y=8x-5, 6y=5-x respectively? 6y=5-x is equivalent to y=-x/6+5/6. so the slopes of the two equations are 8 and -1/6 respectively. apparently 8*(-1/6)≠-1, and therefore they are not perpendicular.